# Ageing-aware Energy Management for Residential Multi-Carrier Energy Systems

### v3

@@ Abstract
In the context of building electrification, the operation of distributed energy resources integrating multiple energy carriers (electricity, heat, mobility) poses a significant challenge due to the nonlinear device dynamics, uncertainty, and computational issues. As such, energy management systems seek to decide the power dispatch in the best way possible. The objective is to minimize and balance operative costs (energy bills or asset degradation) with user requirements (mobility, heating, etc.). Current energy management uses empirical battery ageing models outside of their specific fitting conditions, resulting in inaccuracies and poor performance. Moreover, the link to thermal systems is also overlooked. This paper presents an ageing-aware day-ahead algorithm for electrified buildings that incorporates physics-based battery ageing models. The models distinguish between energy storage systems and make explicit the trade-off between grid cost and battery degradation. The proposed day-ahead algorithm can either cut down on grid costs or extend battery lifetime (electric vehicle or stationary battery packs). Moreover, it exploits the differences between cathode chemistries improving grid costs by 25% when using LFP cells, with respect to NMC cells. Finally, the performance using aged batteries is also enhanced with 35% grid cost observed savings, when passing from new to aged batteries in the summer.
@@
journal: Advances in Applied Energy
@@
Integrated physics-based degradation models into optimal predictive energy management for multi-carrier buildings. This allows users to trade between grid cost reductions and battery lifetime extension.
Novel algorithm can distinguish between cathode chemistries, with LFP cells achieving lower grid costs and capacity fade than their NMC counterpart. Performance is also improved with aged batteries, while traditional benchmarks decrease their grid cost and capacity fade.
@@ 1 Introduction
The decarbonization of the economy as a whole is a significant challenge for modern societies. In particular, the sustainable transformation of both the energy and transport sectors poses significant technical and cultural challenges [1]. Both transitions meet/couple in the population’s homes where electricity, mobility, or heat are needed. Thus, possible synergies between the three systems can be exploited to achieve the desired decarbonization, freedom, resiliency, and cost savings at the local or aggregated level [2]. The successful exploitation of such coupling thus needs to be carefully tailored and built into the design of modern multi-carrier energy systems [3, 2, 4, 5, 6, 7, 8, 9]. This necessarily leads to advanced energy management systems (EMS) that schedule and control the distributed energy resources (DER) [10, 11, 6, 9, 12]. Thus, the EMS needs to handle uncertainty introduced by electric vehicles [13], solar generation and loads, as well as the battery degradation [14, 15, 16, 17, 18]. The main goal of this work is to address this last point.
@@
To dispatch and operate residential multicarrier energy systems (MCES), the literature suggests Model Predictive Control (MPC) [12, 6], stochastic optimization [19], reinforcement learning (RL) [9, 6, 7] and many others. Usually, the basis of such advanced systems is a day-ahead plan or dispatch that schedules the power of the assets along the day [20, 4, 10, 11, 21, 22]. This planner is usually an optimization-based system that uses approximated deterministic forecasts of certain inputs to schedule the different assets. The decisions taken are then implemented and modified in real time. The optimization models have to model the representative aspects of the different assets of the energy system. This includes their power limits, dynamics, and operational constraints. The standard approach is to limit the models to simplified linear or quadratic forms, overlooking most technological particularities [2, 9, 6, 23, 24, 25, 26]. In particular, when several energy storage systems are present, the EMS has to decide which storage system to use and when. In this regard, battery energy storage systems (ESS) are limited by their ageing [15, 16, 17, 18, 27, 5, 10] as well as by their availability for mobility [5, 13]. Current battery energy storage system (BESS) control decisions impact its remaining lifetime, thus, having a control-oriented predictive model that accurately reflects the cost of current decisions is key for optimal operation [28]. The bigger the bias between the optimization model and reality the less trustworthy are the decisions and predictions made.
@@
Unfortunately, although physical ageing mechanisms have been studied and modeled [29, 30, 14, 31, 32, 5], they have been partially incorporated into EMS design of residential MCES through empirical ageing models [5, 10, 23, 20]. On the other hand, the interaction of electricity and heat is becoming more relevant as heating electrification intensifies, and single-carrier optimization might lead to underperformance and cost inefficiencies. In particular, electric vehicle (EV) s and heat pump (HP) s are possible sources of grid congestion [33]. When investigating MCES, Ceusters et al. [6, 7] use first-order linear models for both BESS and thermal energy storage system (TESS), neglecting any differences between their dynamics. Similarly, Ye et al. [9] does not mention any difference between storage systems nor include EV s in their system. Alpizar-Castillo, et al [34] focuses on thermal dynamics, only incorporating BESS with linear models, without including EV s. Other works only focus on the electrical carrier without including coupling electrical heating and storage [4, 10, 16, 35]. These last works apply different battery models to describe their key variables, such as state-of-charge $SoC$, terminal voltage $v_{t}$, and state-of-health $SoH$. In this multi-carrier context, the optimization needs accurate cost functions and predictive models to decide in which carrier to store energy. Since batteries are considerably expensive, optimizing operations to extend their lifetime becomes an attractive option to reduce costs.
@@
Battery-ageing models fall within two categories: empirical or physics-based (PB) [17, 5, 27]. The first are the most widely used in the literature due to their simplicity. They are obtained by performing long standardized calendar and cycling ageing tests [27, 5]. Unfortunately, empirical degradation models only have interpolation capabilities, usually use non-linear equations, represent a limited number of operating conditions (average C-rate, minimum $SoC$, etc.), are prone to overfitting, and are chemistry dependent [32, 36]. On the other hand, PB models are built through first-principles and specialized tests to identify individual degradation mechanisms [29, 31, 27]. They have extrapolation features, can be expressed in the state-space form, account for several cathode chemistries, and represent a wide range of operating conditions. However, they are also non-linear and, in general, non-convex [29, 30, 37, 27, 31]. The integration of PB ageing models into the operation of BESS has been recently studied at the battery management system (BMS) level for standalone and EV applications [14, 38, 39, 15, 16, 40] usually through control-oriented physics-based reduced order model (PBROM). In the cited references, important cost savings were achieved either by preserving battery lifetime or by making an explicit trade-off between the grid costs and capacity fade, even when implementing optimization horizons of a day or less. To the author’s best knowledge, their integration into applications where the BESS interacts with more assets, such as transmission grids, microgrids, industry, and in particular buildings, has not been extensively researched yet [28].
@@
In summary, current optimization-based approaches that do use PBROM to actively trade off between degradation and economic benefits are restricted to standalone battery systems [14, 30, 15, 41, 39, 38]. When the battery is integrated with larger systems (solar photovoltaics (SPV), EV, microgrid, etc.), battery dynamics are usually simplified and empirical models are used in most literature [10, 16, 35]. None of the above references include the integration with a thermal carrier or multiple storage devices, and if they do, they do not include ageing models, [42, 24]. This paper aims to bridge these three gaps.
@@
This paper bridges the previously mentioned gaps by introducing an innovative framework that integrates ageing physics-based reduced-order models (PBROM) directly into the EMS planning algorithm of multi-carrier systems. Current Energy Management Systems (EMS) for residential multi-carrier energy systems (MCES) often plan day-ahead operations without a complete understanding of how their decisions impact the long-term health and lifetime of their storage assets. This integration fundamentally changes how the EMS operates, moving from a short-sighted, one-day optimization to a more holistic, ageing-aware strategy. Our approach enables the system to choose the optimal power dispatch of each storage system based on:
@@ 2 Modeling and Optimal Planner
Our planning algorithm is an optimization-based secondary controller that minimizes energy cost and battery ageing. A schematic of the MCES and the EMS is presented in Fig. 1. The system is composed of SPV, BESS, EV, power electronic interface (PEI), HP, solar thermal (ST), TESS, grid connection and loads. On the left, the EMS decides the day-ahead schedule $P_{a,t}^{*},P_{a,t+1}^{*},\cdots,P_{a,t+24hs}^{*}$, passing it down to the MCES simulator. The MCES simulator feeds back the state measurements $S_{a,t},S_{a,t+1},\cdots,S_{a,t+24hs}$ to continue with the loop.
The following section describes the EMS models, following the Universal Modeling Framework (UMF) by Powell [43, 44, 45]. For a given system size, the objective is to handle the operation cost, which is composed of three parts: the net cost of energy from the grid $C_{\textrm{grid}}$, the degradation cost of losing storage capacity $C_{\textrm{loss}}$, and a penalty for not charging the EV $p_{\textrm{SoCDep}}$. The grid cost and the degradation cost are cumulative objectives because the goal is to optimize them through time, while the penalty for not charging the EV to the desired $SoC$ level is only a point reward at departure times $t_{\textrm{dep}}$. The sequential decision problem (SPD) is then:
@@
$P_{a,t}^{*}=X^{\pi}_{t}(S_{a,t})\in\mathcal{P}$ $\forall a\in\mathbb{A}$ (1c)
$S_{a,t}\in\mathcal{S}$ $\forall a\in\mathbb{A}$ (1d)
@@
$C_{\textrm{grid}}=w_{\text{grid}}\sum_{t=0}^{T}{(\lambda_{\textrm{buy},t}.P^{+}_{\textrm{grid},t}-\lambda_{\textrm{sell},t}.P^{-}_{\textrm{grid},t}).\Delta t}$ (3a)
@@
where $S_{a,t}$ is the state vector, $P^{*}_{a,t}$ is the optimal decision for timestep $t$, $W_{t+1}$ and is an exogenous process that introduces new information after making a decision. The mappings $S_{a,t}^{M}(.)$, and $X^{\pi}_{t}(.)$ are the transition function and optimal policy, respectively. The first is a set of equations describing the states and parameter evolution, and the second is the algorithm that finds the setpoints. The vector $\theta_{a,t}$ contains all the parameters of each asset $a$ and changes over time $t$. The subindex $a\in\mathbb{A}$ corresponds to the assets shown in Fig. 1. The index $b$ denotes the electric storage assets. The evaluation/simulation time window is $T$ and the timestep $\Delta t=15$min. $C_{\textrm{loss}}$ is explained in Section 3.3.2 and the penalty $p_{\textrm{SoCDep}}$ in Section 3.4.
@@
    The order of the subscript is ”name, device, time index”.
@@ 3 Policy design
Figure 2: Deterministic DLA policy as a day-ahead planner.
@@
As mentioned before, the SPD in Eq. 1 is a state-dependent problem where current states influence future decisions. As such, Direct Lookahead (DLA) policies are commonly used in the literature to solve these problems. Two common examples of this policy family are optimal control strategies and stochastic dual dynamic programming. For this work, we focus on day-ahead planning, which is a subset of optimal control where deterministic inputs (forecast medians in this case) are used to decide the actions for the incoming day. The process is shown in Fig. 2. The scheduler takes inputs $\tilde{B}_{a,tt^{\prime}}$ to decide the power dispatch $\tilde{P}^{*}_{a,tt^{\prime}}$ , optimizing over policy time $t^{\prime}\in[t,t+H]$ with $H$ the optimization horizon.
@@
More specifically, a DLA policy based on the devices’ approximated dynamic models. In this way, the EMS plans future actions based on approximate predictive models of devices. Approximation is denoted with $\tilde{}$ . The policy is then solving an approximated economic optimal control problem (OCP) of the form:
$\min_{\tilde{P}^{*}_{a,tt^{\prime}}\in\mathcal{P}}\quad$ $\tilde{C}_{\textrm{grid}}+\tilde{C}_{\textrm{loss}}+\tilde{p}_{\textrm{SoCDep}}+\tilde{p}_{\textrm{TESS}}$ (8a)
s.t. $\tilde{S}_{a,tt^{\prime}+1}=\tilde{S}^{M}_{a,t}\left(\tilde{R}_{a,tt^{\prime}},\tilde{P}^{*}_{a,tt^{\prime}},\tilde{B}_{a,tt^{\prime}}\right)$ (8b)
$\tilde{S}_{a,tt^{\prime}}\in\mathcal{S}$ $\forall a\in\mathbb{A}$ (8c)
where
@@
$\tilde{p}_{\textrm{TESS}}=w_{\textrm{TESS}}.\sum_{t^{\prime}=t}^{H}\max\left(0,\tilde{SoC}_{\textrm{TESS},tt^{\prime}}-\overline{SoC}_{\textrm{TESS}}\right)\Delta t$ (9)
@@
In our policy, the upper limit constraint of the TESS, $\overline{SoC}_{\textrm{TESS}}$, is implemented as a soft constraint to avoid infeasibilities during initialization or feedback. The penalty in the objective steers the $SoC_{\textrm{TESS},t}$ towards the feasible region when the weight $w_{\textrm{TESS}}$ is high enough.
@@
The deterministic optimization problem in Eq. 8 approximates the real stochastic one by using forecasts, stored in $B_{a,t}$, and approximated models for the transition function $\tilde{S}_{a,t}^{M}$. In this model, the time $t$ is the time at which the DLA policy is created and $t^{\prime}$ is the time inside the policy itself. Note the subtle difference between the approximated dynamics $\tilde{S}_{a,t}^{M}$ and the real ones $S_{a,t}^{M}$. This is not to be overlooked because the assumption that the predictions made by the policy $\pi$ hold true can lead to disappointing results in real-world applications. Making these distinctions early in design reveals important insights for future stages. In this work, the energy management algorithm (EMA) has an approximated model $\tilde{S}_{a,t}^{M}$ to decide the setpoints $\tilde{P}^{*}_{a,tt^{\prime}}$ to be implemented in a simulator $S_{a,t}^{M}$ containing detailed fidelity models. In the future, the simulator might as well grow enough to be considered a digital twin of the real building.
@@
$X^{\pi}_{t}(S_{a,t})=\arg\min_{P_{a,t},...,P_{a,t+H}}\tilde{C}_{\textrm{grid}}+\tilde{C}_{\textrm{loss}}+\tilde{p}_{\textrm{SoCDep}}+\tilde{p}_{\textrm{TESS}}$ (10)
@@
A sequence matrix containing power setpoints and states from $t$ to $t+H$ is defined as:
$\mathcal{P}_{a,[t,t+H]}=[P_{a,t},P_{a,t+1},...,P_{a,t+H}]$ (11)
$\mathcal{S}_{a,[t,t+H]}=[S_{a,t},S_{a,t+1},...,S_{a,t+H}]$ (12)
The approximate transition function $\tilde{S}_{a,t}^{M}(.)$ is the compendium of the equations specified in the rest of this section. In the remainder of this section, all equations will be presented just in terms of $t$ for the sake of simplicity. However, the reader must remember that when inside the policy, $\pi$ they are defined under the policy’s time $t^{\prime}$.

### v4

@@ Abstract
In the context of building electrification, the operation of distributed energy resources integrating multiple energy carriers (electricity, heat, mobility) poses a significant challenge due to the nonlinear device dynamics, uncertainty, and computational issues. As such, energy management systems seek to decide the power dispatch in the best way possible. The objective is to minimize and balance operative costs (energy bills or asset degradation) with user requirements (mobility, heating, etc.). Current energy management uses empirical battery ageing models outside of their specific fitting conditions, resulting in inaccuracies and poor performance. Moreover, the link to thermal systems is also overlooked. This paper presents an ageing-aware nonlinear economic model predictive controller for electrified buildings that incorporates physics-based battery ageing models. The models distinguish between energy storage systems (chemistry, ageing state, etc.) and make explicit the trade-off between grid cost and battery degradation. The proposed algorithm can either cut down on grid costs or extend battery lifetime (electric vehicle or stationary battery packs). Additionally, substituting NMC cells with LFP chemistries optimizes grid performance during the summer, yielding a 10% grid cost reduction and a 20% decrease in degradation. Finally, the grid cost and degradation of the presented MPC when using aged batteries are improved with respect to the state of the art by 10% and 5% respectively, in periods with high solar generation and low thermal loads like summer.
@@
journal: Journal of Energy Storage
@@
Integrated physics-based degradation models into optimal predictive energy management for multi-carrier buildings. This allows users to trade-off between grid cost reductions and battery lifetime extension.
Novel algorithm can distinguish between cathode chemistries, with LFP cells achieving lower grid costs and capacity fade than their NMC counterpart.
Integrating physics-based ageing improves EMS performance with aged batteries, while traditional benchmarks increase their grid cost and capacity fade.
@@ 1 Introduction
The decarbonization of the economy as a whole is a significant challenge for modern societies. In particular, the sustainable transformation of both the energy and transport sectors poses significant technical and cultural challenges [IEA2021]. Both transitions couple in the population’s homes where electricity, mobility, or heat are needed. Thus, possible synergies between the three systems can be exploited to achieve the desired decarbonization, freedom, resiliency, and cost savings at the local or aggregated level [Geidl2007]. The successful exploitation of such coupling thus needs to be carefully tailored and built into the design of modern multi-carrier energy systems [Andersson2005, Geidl2007, Vermeer2020, Vermeer2022, Ceusters2021, Ceusters2023, VanDerMeer2018, Ye2020]. This necessarily leads to advanced energy management systems (EMS) that schedule and control the distributed energy resources (DER) [Vermeer2022b, Vermeer2022AData, Ceusters2021, Ye2020, EsmaeelNezhad2022]. Thus, the EMS needs to handle uncertainty introduced by electric vehicles [Alexeenko2023AchievingStudy], solar generation and loads, as well as the battery degradation [Xavier2021, Jin2022, Li2019, Plett2015, Plett2016]. The main goal of this work is to address this last point.
@@
To dispatch and operate residential multicarrier energy systems (MCES), the literature suggests Model Predictive Control (MPC) [EsmaeelNezhad2022, Ceusters2021], stochastic optimization [Chen2012], reinforcement learning (RL) [Ye2020, Ceusters2021, Ceusters2023] and many others. Usually, the basis of such advanced systems is a day-ahead plan or dispatch that schedules the power of the assets along the day [Risbeck2018Mixed-IntegerSystems, Vermeer2020, Vermeer2022b, Vermeer2022AData, Jouini2024PredictiveStudy, Li2021IntradaySystem]. This planner is usually an optimization-based system that uses approximated deterministic forecasts of certain inputs to schedule the different assets. The decisions taken are then implemented and modified in real time. The optimization models have to model the representative aspects of the different assets of the energy system. This includes their power limits, dynamics, and operational constraints. The standard approach is to limit the models to simplified linear or quadratic forms, overlooking most technological particularities [Geidl2007, Ye2020, Ceusters2021, Mariano-Hernandez2021ADiagnosis, Yang2023BuildingPricing, Mittelviefhaus2022ElectrificationMitigation, Su2025DynamicSystems, Karthikeyan2025EnhancingSystems]. This nonlinearities can be the efficiency of a heat-pump, the voltage response of a battery or the heat losses of a building. In particular, when several energy storage systems are present, the EMS has to decide which storage system to use and when. In this regard, battery energy storage systems (ESS) are limited by their ageing [Jin2022, Li2019, Plett2015, Plett2016, Reniers2019, Vermeer2022, Vermeer2022b] or their availability for mobility [Vermeer2022, Alexeenko2023AchievingStudy]. Current battery energy storage system (BESS) control decisions impact its remaining lifetime, thus, having a control-oriented predictive model that accurately reflects the cost of current decisions is key for optimal operation [VegaGarita2025TheReview]. The bigger the bias between the optimization model and reality the less trustworthy are the decisions and predictions made.
@@
On the other hand, the interaction of electricity and heat is becoming more relevant as heating electrification intensifies, and single-carrier optimization might lead to under performance and cost inefficiencies. When investigating MCES, Ceusters et al. [Ceusters2021, Ceusters2023] use first-order linear models for both BESS and thermal energy storage system (TESS), neglecting any differences between their dynamics. Similarly, Ye et al. [Ye2020] does not mention any difference between storage systems nor include electric vehicle (EV) s in their system. Alpizar-Castillo, et al [Alpizar-Castillo2024ModellingHouse] focuses on thermal dynamics, only incorporating BESS with linear models, without including EV s. Other works only focus on the electrical carrier without including coupling electrical heating and storage [Vermeer2020, Vermeer2022b, Li2019, Li2023EnsembleManagement]. These last works apply different battery models to describe their key variables, such as state-of-charge $SoC$, terminal voltage $v_{t}$, and state-of-health $SoH$. In this multi-carrier context, the optimization needs accurate cost functions and predictive models to decide in which carrier to store energy. If no dynamic distinction is made between ESS then the predicted operation will be far from reality leading to over investment on inefficient technologies. Thus, this work integrates electrical and thermal storage models together in the same optimization, but using different dynamic models to reflect their particular technologies accordingly.
@@
Unfortunately, although physical ageing mechanisms have been studied and modeled [Okane2022, Jin2017, Xavier2021, Prada2013Simulations, Wang2014, Vermeer2022], they have been partially incorporated into EMS design of residential MCES through empirical ageing models [Vermeer2022, Vermeer2022b, Mariano-Hernandez2021ADiagnosis, Risbeck2018Mixed-IntegerSystems]. Battery-ageing models fall within two categories: empirical or physics-based (PB) [Plett2015, Vermeer2022, Reniers2019]. The first are the most widely used in the literature due to their simplicity. They are obtained by performing long standardized calendar and cycling ageing tests [Reniers2019, Vermeer2022]. Unfortunately, empirical degradation models only have interpolation capabilities, usually use non-linear equations, can only predict regular cycles (average C-rate, minimum $SoC$, etc.), are prone to overfitting, and are chemistry dependent [Wang2014, Schmalstieg2014, Reniers2019]. On the other hand, PB models are built through first-principles and specialized tests to identify individual degradation mechanisms [Okane2022, Prada2013Simulations, Reniers2019]. They have extrapolation features, can be expressed in the state-space form, account for several cathode chemistries, and represent dynamic profiles such as EV s, utility scale BESS or any irregular load. However, they are also non-linear and, in general, non-convex [Okane2022, Jin2017, Purewal2014DegradationModel, Reniers2019, Prada2013Simulations]. The integration of PB ageing models into the operation of BESS has been recently studied at the battery management system (BMS) level for standalone and EV applications [Xavier2021, Reniers2023, Reniers2021, Jin2022, Li2019, Li2024NonlinearCharging] usually through control-oriented physics-based reduced order model (PBROM). In the cited references, important cost savings were achieved either by preserving battery lifetime or by making an explicit trade-off between the grid costs and capacity fade, even when implementing optimization horizons of a day or less. To the author’s best knowledge, their integration into applications where the BESS interacts with more assets, such as transmission grids, microgrids, industry, and in particular buildings, has not been extensively researched yet [VegaGarita2025TheReview]. In this work, physics-based models unlock a new level of EMS intelligence: the ability to differentiate unique battery chemistries and evolve alongside the cell’s actual aging process, ensuring optimal dispatch across the entire lifecycle of the multi-carrier system.
Table 1: Summary of Literature Review.
Loads / Carriers Flexibility Battery modelling Ref. Application Elec. Load Space Heating Nat.Gas TESS EV BESS Heat Pump Performance Degradation Dynamic Profile Cathode Chemistry Used and New Cells [Ceusters2021, Ceusters2023, Ceusters2023AnSystems] Energy Hub Coulomb counting [Vermeer2022b] HEMS Flow-based Coulomb counting Empirical [Geidl2007] Energy Hub Coulomb counting [Ye2020] HEMS Coulomb counting [Yang2023BuildingPricing] HEMS Coulomb counting [Karthikeyan2025EnhancingSystems] Energy Hub Coulomb counting Empirical [Su2025DynamicSystems] Energy Hub Flow-based Coulomb counting Empirical [Reniers2021, Reniers2018] Utility scale PBROM * * [Xavier2021] Fast-charging PBROM * * [Jin2022] Fast-charging ECM PBROM * [Li2019] HEMS PB ECM * * [Li2023EnsembleManagement] HEMS PB ECM * * [Dorronsoro2025BatteryModels] HEMS ECM Empirical [Park2023OptimalApproach] Microgrid Coulomb Counting Empirical This work HEMS ECM PBROM Note *: Capable, but no Case Study
@@
The most relevant works can be found in Table 1. In summary, current optimization-based approaches that do use PBROM to actively trade off between degradation and economic benefits are restricted to standalone battery systems [Xavier2021, Jin2017, Jin2022, Reniers2018, Reniers2021, Reniers2023, Li2024NonlinearCharging]. When the battery is integrated with larger systems (solar photovoltaics (SPV), EV, microgrid, etc.), battery dynamics are usually simplified and empirical models are used in most literature, leading to inaccurate degradation predictions and under utilization [Vermeer2022b, Li2019, Li2023EnsembleManagement, Dorronsoro2025BatteryModels, Park2023OptimalApproach]. None of the above references include the integration with a thermal carrier or multiple storage devices, and if they do, they do not include ageing models, [Su2025DynamicSystems, Yang2023BuildingPricing, Geidl2007, Ceusters2023, Karthikeyan2025EnhancingSystems].
@@
This paper bridges the previously mentioned gaps by introducing an innovative economic MPC framework that integrates ageing physics-based reduced-order models (PBROM) directly into the EMS algorithm of multi-carrier systems. Current Energy Management Systems (EMS) for residential multi-carrier energy systems (MCES) often operate without a complete understanding of how their decisions impact the long-term health and lifetime of their storage assets. This integration fundamentally changes how the EMS operates, moving from a short-sighted, price-based optimization to a more holistic, ageing-aware strategy. Our approach enables the system to choose the optimal power dispatch of each storage system based on:
@@ 2 Modeling and Optimal Planner
Our EMS is an optimization-based secondary controller that minimizes energy cost and battery ageing. A schematic of the MCES and the EMS is presented in Fig. 1. The system is composed of SPV, BESS, EV, power electronic interface (PEI), heat pump (HP), solar thermal (ST), TESS, grid connection and loads. On the left, the EMS decides the power dispatch $P_{a,t}^{*}$ at $t\in\mathcal{D}_{t}$ where $\mathcal{D}_{t}=[0,T]$, passing it down to the MCES simulator. The MCES simulator feeds back the state measurements $S_{a,t}$ to continue with the loop.
The following section describes the EMS models, following the Universal Modeling Framework (UMF) by Powell [Powell2022]. For a given system size, the objective is to handle the operation cost, which is composed of three parts: the net cost of energy from the grid $C_{\textrm{grid}}$, the degradation cost of losing storage capacity $C_{\textrm{loss}}$, and a penalty for not charging the EV $p_{\textrm{SoCDep}}$. The grid cost and the degradation cost are cumulative objectives because the goal is to optimize them through time, while the penalty for not charging the EV to the desired $SoC$ level is only a point reward at departure times $t_{\textrm{dep}}$. The sequential decision problem (SPD) is then:
@@
$P_{a,t}^{*}=X^{\pi}_{t}(S_{a,t})\in\mathcal{P}$ $\forall a\in\mathbb{A}$ $,\ t\in\mathcal{D}_{t}$ (1c)
$S_{a,t}\in\mathcal{S}$ $\forall a\in\mathbb{A}$ $,\ t\in\mathcal{D}_{t}$ (1d)
@@
$C_{\textrm{grid}}=w_{\text{grid}}\sum_{t=0}^{T}{\left(\lambda_{\textrm{buy},t}.P^{+}_{\textrm{grid},t}-\lambda_{\textrm{sell},t}.P^{-}_{\textrm{grid},t}\right).\Delta t}$ (3a)
@@
where $S_{a,t}$ is the state vector, $P^{*}_{a,t}$ is the optimal decision for timestep $t$, $W_{t+1}$ and is an exogenous process that introduces new information after making a decision. The mappings $S_{a,t}^{M}(.)$, and $X^{\pi}_{t}(.)$ are the transition function and optimal policy, respectively. The first is a set of equations describing the states and parameter evolution, and the second is the algorithm that finds the setpoints. The vector $\theta_{a,t}$ contains all the parameters of each asset $a$ and changes over time $t$. The subindex $a\in\mathbb{A}$ corresponds to the assets shown in Fig. 1. The index $b$ denotes the electric storage assets. The evaluation/simulation time window is $\mathcal{D}_{t}\in[0,T]$ and the timestep $\Delta t=15$min. $C_{\textrm{loss}}$ is explained in Section 3.3.2 and the penalty $p_{\textrm{SoCDep}}$ in Section 3.4.
@@
    The order of the subscript is "name, device, time index".
@@ 3 Policy design
Figure 2: Deterministic DLA policy with a day-ahead planner and economic MPC layer.
@@
As mentioned before, the SPD in Eq. 1 is a state-dependent problem where current states influence future decisions. As such, Direct Lookahead (DLA) policies are commonly used in the literature to solve these problems. Two common examples of this policy family are optimal control strategies and stochastic dual dynamic programming. For this work, we focus on economic nonlinear MPC [Grune2017NonlinearAlgorithms], which is a subset of optimal control where deterministic inputs (forecast medians in this case) are used to decide the actions for the incoming day. The process is shown in Fig. 2. The policy $X^{\pi}$ takes inputs $\tilde{B}_{a,tt^{\prime}}$ to decide the power dispatch $\tilde{P}^{*}_{a,tt^{\prime}}$ , optimizing over policy time $t^{\prime}\in[t,t+H^{\text{DA/MPC}}]$ with $H^{\text{DA/MPC}}$ the optimization horizons.
@@
More specifically, a DLA policy based on the devices’ approximated dynamic models $\tilde{S}_{t}^{M}(\cdot)$. In this way, the EMS plans future actions based on approximate predictive models of devices. Approximation is denoted with $\tilde{}$ . The policy is then solving the implicit economic Model Predictive Control (eMPC) given by:
$\min_{\tilde{P}^{\text{DA/MPC}}_{a,tt^{\prime}}\in\mathcal{P}}\quad$ $\tilde{J}^{\text{DA/MPC}}$ (8a)
s.t. $\tilde{S}_{a,tt^{\prime}+1}^{\text{DA/MPC}}=\tilde{S}^{M}_{a,t}\left(\tilde{R}_{a,tt^{\prime}}^{\text{DA/MPC}},\tilde{P}^{\text{DA/MPC}}_{a,tt^{\prime}},\tilde{B}^{\text{DA/MPC}}_{a,tt^{\prime}}\right)$ (8b)
$\tilde{S}_{a,tt^{\prime}}^{\text{DA/MPC}}\in\mathcal{S}$ $\forall a\in\mathbb{A}$ (8c)
where:
$\tilde{J}^{\text{DA/MPC}}=\tilde{C}_{\textrm{grid}}^{\text{DA/MPC}}+\tilde{C}_{\textrm{loss}}^{\text{DA/MPC}}+\tilde{p}_{\textrm{SoCDep}}^{\text{DA/MPC}}+\tilde{p}_{\textrm{TESS}}^{\text{DA/MPC}}$ (9a)
@@
$\tilde{p}_{\textrm{TESS}}^{\text{DA/MPC}}=w_{\textrm{TESS}}.\sum_{t^{\prime}=t}^{H}\max\left(0,\tilde{SoC}^{\text{DA/MPC}}_{\textrm{TESS},tt^{\prime}}-\overline{SoC}_{\textrm{TESS}}\right)\Delta t$ (9b)
@@
In our policy, the upper limit constraint of the TESS, $\overline{SoC}_{\textrm{TESS}}$, is implemented as a soft constraint to avoid infeasibilities during initialization or feedback. The penalty in the objective steers the $SoC_{\textrm{TESS},t}$ towards the feasible region when the weight $w_{\textrm{TESS}}$ is high enough. The eMPC is an adaptation of the policy presented in [Slaifstein2026SequentialMPC]. The only difference between the two layers (DA and MPC) is the length of the horizon and the terminal conditions, which will be analyzed in the Section 3.3.
@@
The deterministic optimization problem in Eq. 8 approximates the real stochastic one by using forecasts, stored in $\tilde{B}_{a,tt^{\prime}}$, and approximated models for the transition function $\tilde{S}_{a,t}^{M}$. In this model, the time $t$ is the time at which the DLA policy is created and $t^{\prime}$ is the time inside the policy itself. Note the subtle difference between the approximated dynamics $\tilde{S}_{a,t}^{M}$ and the real ones $S_{a,t}^{M}$. This is not to be overlooked because the assumption that the predictions made by the policy $\pi$ hold true can lead to disappointing results in real-world applications. Making these distinctions early in design reveals important insights for future stages. In this work, the energy management algorithm (EMA) has an approximated model $\tilde{S}_{a,t}^{M}$ to decide the setpoints $\tilde{P}^{*}_{a,tt^{\prime}}$ to be implemented in a simulator $S_{a,t}^{M}$ containing detailed fidelity models. In the future, the simulator might as well grow enough to be considered a digital twin of the real building.
@@
$X^{\textrm{DA/MPC}}_{t}(S_{a,t})=\arg\min_{P_{a,t}^{\textrm{DA/MPC}}}\tilde{J}^{\textrm{DA/MPC}}$ (10)
@@
The approximate transition function $\tilde{S}_{a,t}^{M}(.)$ is the compendium of the equations specified in the rest of this section. In the remainder of this section, all equations will be presented just in terms of $t$ for the sake of simplicity. However, the reader must remember that when inside the policy $X^{\pi}$ they are defined under the policy’s time $t^{\prime}$.